OpenStudy (anonymous):

Zach borrowed four thousand dollars from his grandfather to buy a used car so he could travel to and from work. After working for 2 years, he has saved four thousand seven hundred twenty dollars to pay back to his grandfather. At what interest rate did Zach borrow the money?

OpenStudy (anonymous):

eight percent nine percent fifteen percent forty-two percent

undeadknight26 (undeadknight26):

one sec.

undeadknight26 (undeadknight26):

Google this question open study has it alot...my computer isn't loading sorry...

OpenStudy (anonymous):

k thanks any way

undeadknight26 (undeadknight26):

sorry man!

OpenStudy (anonymous):

I = PRT I/PT = R 720/(4000)(2) = R 720/8000 = r .09 = r 9% = r

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem :)